# Mercury: A Computational Efficiency Benchmark for Code Large Language Models

### v2

@@ Abstract
Amidst the recent strides in evaluating Large Language Models for Code (Code-LLMs), existing benchmarks have mainly focused on functional correctness, overlooking the importance of computational efficiency. To fill the gap, we present Mercury, the first computational efficiency benchmark for Code-LLMs. It comprises 1,889 Python tasks, each with adequate solutions to support a runtime distribution. Based on the distribution, we introduce a new metric $Beyond$, which computes a runtime-percentile-weighted $Pass$ score to reflect functional correctness and computational efficiency simultaneously. On Mercury, leading Code-LLMs can achieve 67% on $Pass$, while less then 50% on $Beyond$. Given that an ideal $Beyond$ score would be aligned with the $Pass$ score, it indicates that while Code-LLMs exhibit impressive capabilities in generating functionally correct code, there remains a notable gap in their efficiency. Finally, our empirical experiments reveal that Direct Preference Optimization (DPO) serves as a robust baseline for enhancing computational efficiency compared with Supervised Fine Tuning (SFT), which paves a promising avenue for future exploration of efficient code generation. [^1] Our code and data are available on GitHub: https://github.com/Elfsong/Mercury..
@@ 1 Introduction
The domain of code generation, which aims to empower computers to autonomously generate code based on natural language task instructions (NL2Code), constitutes a longstanding and fascinating research objective [42]. The recent emergence of Large Language Models (LLMs) has spurred the development of a new wave of code generation models [40, 31, 35, 11, 27, 4], which leverage the impressive language understanding and generative capabilities of LLMs to drive forward the ambitious goal of producing high-quality and machine-generated code.
In parallel with the model development, recent code generation benchmarks mainly focus on evaluating functional correctness in the fuzz testing style [25]. This testing approach assesses the outcome congruence between the generated and anticipated solutions by executing bespoken test cases. For instance, HumanEval [8] and MBPP [3] collected a small but high-quality set of handcrafted tasks, and EvalPlus [24] further consolidates these two above benchmarks by augmenting the test scope. On the contrary, APPS [14] widely gathered over 5,000 public coding tasks from online platforms. Despite these strides, there is a discernible oversight in current code generation benchmarks concerning the computational efficiency evaluation, although that is critical in real-world software development [43, 44]. Moreover, for most tasks in the above benchmarks, handcrafting diverse solutions and comprehensive test cases to cover all possible scenarios is a time-consuming endeavor [24]. In light of these findings, we highlight several vital limitations inherent in the existing code generation benchmarks:
Figure 1: While both LLM-generated codes successfully follow the task instruction, the right snippet notably excels in computational efficiency, executing in a mere 121 ms compared to the 5,714 ms consumed by the left snippet. On the premise of ensuring functional correctness, optimizing computing efficiency is a ubiquitous objective in software development.
@@
    Absence of Computational Efficiency Evaluation. Existing code generation benchmarks focus on assessing functional correctness while overlooking the evaluation of computational efficiency [8, 3, 14]. As illustrated in Figure 1, despite both code snippets can handle the sorting task functionally, the right efficient solution (121 ms) is nearly 50 times faster than the left inefficient solution (5,714 ms). This striking runtime differentiation underscores the necessity of incorporating computational efficiency assessments within code generation benchmarks to ensure that synthesized code is not only correct but also efficient.
@@
    Insufficient Test Case Coverage. Most code generation benchmarks obtain a limited set of handcrafted test cases for assessing functional correctness, potentially overlooking edge cases and nuanced code behaviors [8, 3]. For example, Figure 3 displays that HumanEval #55 contains only 3 test cases, testing up to the 12th Fibonacci number [8]. Its given canonical solution will quickly reach the recursion depth limitation when computing a larger Fibonacci number (the recursion limitation depends on the environment). Therefore, notwithstanding the generated code satisfies all test cases, such success does not necessarily equate to assurance of functional correctness and much less to computational efficiency.
@@
    Lack of Task Diversity. Another noticeable deficit of existing code generation benchmarks is the insufficient diversity and complexity in their constituent tasks [8, 3, 25]. Since most benchmarks only consist of elementary-level problems, recent Code-LLMs can effortlessly tackle most tasks regardless of their actual capacities [44]. This flaw results in these benchmarks failing to pose a substantial challenge to Code-LLMs and truly reflect their underlying potential.
Computational Efficiency.
@@
Computational efficiency refers to the performance measure of time and space complexity to accomplish a specific task. Efficient code can improve user experience, save energy, and make applications more sustainable and cost-effective. Compared with the scalable memory space, execution time is the performance bottleneck of most programs. Consequently, this work focuses on the time dimension of computational efficiency.
@@
Figure 2: An overview of Mercury dataset. Each Mercury task has a task description, a test case generator, a prompt & entry point, and corresponding solutions. To evaluate computational efficiency, we introduce the $Beyond$ metric, which signifies the runtime percentile of the LLM-generated code on the runtime distribution supported by corresponding solutions. In this example, the LLM-generated code executes in 521 ms, outpacing 86.18% of collected solutions on the runtime distribution. Consequently, the $Beyond$ metric in this case is 86.18%.
Our Benchmark.
@@
In this work, we introduce Mercury, a novel code generation benchmark designed to assess computational efficiency. As depicted in Figure 2, Mercury comprises 1,889 Python programming tasks with three difficulty stratification, which is divided into two datasets for model fine-tuning and evaluation separately. For each task, we assign a test case generator to remedy the shortfall of test case coverage. In measuring computational efficiency, the primary challenge stems from normalizing the absolute runtime across tasks that have diverse runtime ranges. Thus we collect and locally execute numerous solutions for each task to form a runtime distribution and leverage the runtime percentile of LLM-generated code on the distribution instead of the absolute runtime to evaluate computational efficiency. Furthermore, to mitigate performance discrepancies attributed to irrelevant processes and diverse hardware configurations, we set up an isolated sandbox environment for task execution to establish local runtime distributions.
Contribution.
@@
Our work aimed to fill the computational efficiency evaluation gap in code generation benchmarks with the following key contributions:
@@
  • Dataset. We introduce a novel code generation dataset Mercury designed to assess and improve computational efficiency in Section 2, accompanied by an extensible open-source data collection framework for enriching Mercury with more tasks and languages.
@@
  • Metric. We propose the first efficiency-focused code generation metric $Beyond$ and establish a benchmark to evaluate leading Code-LLMs using this metric in Section 3.
@@
  • Baselines. In Section 4, we detail our extensive analysis of baselines to enhance computational efficiency while maintaining functional correctness. Experiment results reveal that despite Code-LLMs excelling in functional correctness, there is still considerable potential to elevate efficiency.
@@
Table 1: A comparison of Mercury to existing NL2Code benchmarks. Mercury distinguishes itself by including a set of distilled high-quality solutions and a dedicated test case generator for each task.
@@
Mercury 256 Online + Filters $\infty$ 18.4 3
Figure 3: An example of insufficient test cases. Even though the code passed all test cases in the dashed-line box, it remains vulnerable to timeout or stack overflow when subjected to a larger input.
@@ 2 Mercury Datasets
We initiate this work by collecting public programming tasks on Leetcode [^2] https://leetcode.com/problemset/algorithms/. Subjecting these questions to a series of filters, we distilled them down to 1,889 high-quality tasks. A difficulty-balanced subset of 256 tasks was randomly selected to form the Mercury-eval benchmark, which obtains an average of 18.4 solutions for each problem. The remaining tasks have been designated as the Mercury-train dataset for baseline training (detailed data distribution is listed in Appendix Table 6). To enhance clarity within this paper, we employ Mercury to denote Mercury-eval unless otherwise specified.
Data Schema.
As illustrated in Figure 2, Mercury offers a unified data schema to streamline the evaluation procedure and bolster further development endeavors. The data scheme encompasses these principal components: (1) Task Description contains the task instruction interpreted into a plain text format, along with illustrative examples and constraints of inputs and outputs. (2) Test Case Generator refers to a Python code snippet designed to automatically produce a comprehensive set of test cases in accordance with the specifications laid out in the task description. (3) Solutions are sampled from Leetcode historical submissions. Each solution within Mercury has undergone rigorous testing, and Locality-Sensitive Hashing [16] is employed to prevent the inclusion of any identical solutions. (4) Prompts and Entry Points where prompts act as the initiating prefixes for LLM code generation and entry points denote the start point for code execution. We delineate the definition of Mercury fields in the Appendix Table 5.
Task Filters.
Mercury tasks originate from public programming problems on Leetcode. To assure the quality and uniformity of the dataset, we distilled gathered tasks based on the following conditions:
@@
Restricted Data Structure. Above the inherent Python data types, Mercury also incorporates two custom data types: Binary Tree and Linked List (the specific structure definitions can be found in Figure 4), which increases Mercury’s diversity and escalates its difficulty level. Tasks that contain other data structures will be removed.
@@
Unique Outputs. Certain Leetcode tasks may permit non-unique answers. For example, a result list can be returned in any order. Evaluating all possible answers can drastically complicate the test case verification process. To eliminate this problem, we harness the corresponding test case generator to generate $N$ test cases $T_{i}=\langle Input_{i},Output_{i}\rangle\ s.t.\ i\in\{0,1,\cdots,N\}$ and execute $T$ on different solutions $S_{m}\ s.t.\ m\in\{0,1,\cdots,M\}$ to observe if all outputs $Output_{i}=S_{m}(Input_{i})$ remain identical. Any tasks that potentially yield non-unique answers were subsequently excluded.
Task Difficulty.
@@
Most existing NL2Code benchmarks predominantly comprise simplistic tasks, leading to a situation where LLMs of varied capabilities address most tasks effortlessly and yield indistinguishable high scores [44, 17]. To alleviate this issue, Mercury inherits the difficulty categorization from Leetcode, i.e., Easy, Medium, and Hard. The stratification aims to probe the upper bounds of Code-LLM capabilities, delivering a more evident distinction between various Code-LLMs.
Test Case Generator.
@@
Manual creation of test cases can be a laborious process. To gather sufficient test cases to conduct an exhaustive assessment, we assign a test case generator for each task, which can produce a full range of test cases to thoroughly evaluate the functional correctness and efficiency of given solutions. Specifically, We reformulate the task description as plain text (see $pretty\_content$ in Table 5) and then feed it into GPT-4 [31] to generate an initial test case generator snippet. To confirm the effectiveness of the initial generator, we subsequently create a batch of test cases (24 cases in our setting) by the generator and submit them to the Leetcode online judge system. Should any of the generated test cases not pass the Leetcode validation, we manually revise the generator until all generated cases can be successfully validated.
3 Computational Efficiency Metric
@@
In the domain of software development, computational efficiency can be defined as the absolute code runtime for executing a given test case set [7]. Nonetheless, a primary obstacle in benchmarking computational efficiency is normalizing runtime measurements across disparate environments. For instance, a sub-optimal solution might have a faster absolute runtime on high-performance hardware than an optimal solution on low-performance hardware. Moreover, different operation systems and code interpreters may also fluctuate the code runtime. Therefore, absolute runtime fails as a consistent and reliable computational efficiency benchmark metric. To address this issue, an intuitive approach involves modeling a devoted runtime distribution for each task and calculating the average runtime percentiles of LLM solution samples over the runtime distribution. With this idea in mind, we proposed a normalized computational efficiency metric $Beyond$:
@@
Where $N$ is the total number of tasks, and $K$ denotes the size of LLM solution samples. For a specific task $n\in N$, $R^{n}$ is the runtime array corresponding to the collected solutions, and $r_{k}^{n}$ denotes the runtime for the $k$-th LLM solution. Runtime is defined as the period from the solution instantiation to the evaluation across all test cases, culminating with a successful termination (More engineering details can be found in Appendix Figure 5). Since any case failure of the $k$-th solution results in $r_{k}^{n}\rightarrow+\infty$ and then $p_{k}^{n}=0$, $Beyond$ reflects functional correctness as well.
Untrusted Code Execution.
Since most Code-LLMs are trained on an extensive code corpus from unverified sources, there is an intrinsic risk that these models may produce malicious code when driven by specific meticulous prompts [8]. The direct execution of synthesized code raises significant security concerns. To alleviate the risk of running untrusted code, we engage a robust sandbox to execute code in an isolated environment. Sandbox details are deliberated in Appendix A.3
Hardware-agnostic Evaluation.
To ensure fair comparison across diverse hardware configurations, we run each task $n$ with corresponding test cases locally and aggregate their runtimes into the runtime array $R^{n}$. Appendix Figure 8 illustrates the $Beyond$ score of two LLMs (‘deepseek-coder-33b’ and ‘deepseek-coder-6.7b’) over three distinct hardware specifications: the micro-tier (0.25 CPU cores), the small-tier (0.5 CPU cores), and the standard-tier (1 CPU core). The results demonstrate that Beyond remains consistent over different hardware configurations.

### v3

# Mercury: A Code Efficiency Benchmark for Code Large Language Models

@@ Abstract
Amidst the recent strides in evaluating Large Language Models for Code (Code LLMs), existing benchmarks have mainly focused on the functional correctness of generated code, neglecting the importance of their computational efficiency. To fill the gap, we present Mercury, the first code efficiency benchmark for Code LLMs. It comprises 1,889 Python tasks, each accompanied by adequate solutions that serve as real-world efficiency baselines, enabling a comprehensive analysis of the runtime distribution. Based on the distribution, we introduce a new metric Beyond, which computes a runtime-percentile-weighted Pass score to reflect functional correctness and code efficiency simultaneously. On Mercury, leading Code LLMs can achieve 65% on Pass, while less than 50% on Beyond. Given that an ideal Beyond score would be aligned with the Pass score, it indicates that while Code LLMs exhibit impressive capabilities in generating functionally correct code, there remains a notable gap in their efficiency. Finally, our empirical experiments reveal that Direct Preference Optimization (DPO) serves as a robust baseline for enhancing code efficiency compared with Supervised Fine Tuning (SFT), which paves a promising avenue for future exploration of efficient code generation.  [^1] Our code and data are available on GitHub: https://github.com/Elfsong/Mercury.
@@ 1 Introduction
The domain of code generation, which aims to empower computers to autonomously generate code based on natural language task descriptions (NL2Code), has long been considered a promising way to facilitate interaction between humans and computers [46, 44]. The recent emergence of Large Language Models (LLMs) has spurred a new wave of NL2Code models [34, 38, 12, 29, 4], which leverage the impressive language understanding and generative capabilities of LLMs to drive forward the ambitious goal of synthesizing high-quality code from natural language instructions.
To measure the quality of code, recent code generation benchmarks mainly focus on evaluating their functional correctness via test case fuzzing [27]. This approach assesses the outcome congruence between the LLM-generated and canonical solutions by executing bespoken test cases. For instance, HumanEval [8] and MBPP [3] collected a small but fine set of handcrafted tasks with test cases. EvalPlus [26] further consolidates these two above benchmarks by augmenting the case scope. On the contrary, APPS [15] widely gathered over 5,000 public coding tasks from online platforms. Despite these strides, there is a discernible oversight in current code generation benchmarks concerning the code efficiency evaluation, although that is critical in software development [45, 47]. Moreover, handcrafting diverse solutions and test cases to cover all scenarios is infeasible [26]. In light of these findings, we highlight vital limitations inherent in the existing code generation benchmarks:
Figure 1: Executing these two LLM-generated codes on 100 test cases. While both codes successfully follow the task instruction and pass all test cases, the right snippet notably excels in code efficiency, completing in a mere 121 ms compared to the 5,714 ms consumed by the left snippet. As Code LLMs become widely used in the real world, code efficiency determines factual productivity, where Mercury can gauge the vital metric.
@@
    Absence of code Efficiency Evaluation. Existing code generation benchmarks focus on assessing functional correctness while overlooking the evaluation of code efficiency [8, 3, 15]. As illustrated in Figure 1, despite both code snippets can handle the sorting task functionally, the right efficient solution (121 ms) is nearly 50 times faster than the left inefficient solution (5,714 ms). This striking runtime differentiation underscores the necessity of incorporating code efficiency assessments within code generation benchmarks, encouraging Code LLMs to produce not only correct but also efficient code.
@@
    Insufficient Test Case Coverage. As shown in Table 1, most code generation benchmarks manually build a small number of test cases or extract the accompanying test cases from existing resources, potentially overlooking edge cases and nuanced code behaviors [8, 3]. For example, Figure 8 displays that HumanEval #55 contains only 3 test cases, testing up to the 12th Fibonacci number [8]. Its given canonical solution will quickly reach the recursion depth limitation when computing a larger Fibonacci number (the recursion limitation depends on the environment). Therefore, notwithstanding the generated code satisfies all test cases, such success does not necessarily equate to assurance of functional correctness and much less to code efficiency.
@@
    Lack of Task Diversity. Another noticeable deficit of existing code generation benchmarks is the insufficient diversity and complexity in their tasks [8, 3, 27]. Since most benchmarks only consist of elementary-level programming tasks, recent Code LLMs can effortlessly tackle most tasks regardless of their actual capacities [47]. This flaw results in these benchmarks failing to pose a substantial challenge to Code LLMs and truly reflect their underlying potential.
@@
Code Efficiency. Code efficiency refers to the performance measure of time and space complexity to accomplish a specific task. Efficient code can improve user experience, save energy, and make applications more sustainable and cost-effective. Compared with the scalable memory space, execution time is the performance bottleneck of most codes. Consequently, this work focuses on the time dimension of code efficiency.
@@
Figure 2: An overview of Mercury dataset. Each Mercury task has a task description, a test case generator, a prompt & entry point, and corresponding solutions. To evaluate code efficiency, we introduce the Beyond metric, which signifies the runtime percentile of the LLM-generated code on the runtime distribution supported by corresponding solutions. In this example, the LLM-generated code executes in 521 ms, outpacing 86.18% of collected solutions on the runtime distribution. Consequently, the Beyond metric in this case is 86.18%.
@@
Our Benchmark. In this work, we introduce Mercury, a novel code generation benchmark designed to assess and improve the code efficiency of Code LLMs. As depicted in Figure 2, Mercury comprises 1,889 Python programming tasks with three difficulty stratification, which is divided into two datasets for model evaluation and fine-tuning separately. For each evaluation task, we assign a test case generator to remedy the shortfall of test case coverage. In measuring code efficiency, the primary challenge stems from normalizing the absolute runtime across tasks that have diverse runtime ranges. Thus, we collect and locally execute numerous historical solutions for each task to form a runtime distribution and leverage the runtime percentile of LLM-generated code on the distribution instead of the absolute runtime to evaluate code efficiency. Furthermore, to mitigate performance discrepancies attributed to irrelevant processes and diverse hardware configurations, we set up an isolated sandbox environment for task execution to establish local runtime distributions.
@@
Contribution. Our work aimed to fill the code efficiency evaluation gap in code generation benchmarks with the following key contributions:
@@
  • Dataset. We collect a novel code generation dataset Mercury designed to assess and improve Code LLM code efficiency in Section 2, accompanied by an extensible open-source data collection framework for enriching Mercury with more tasks and programming languages.
@@
  • Metric. We propose the first efficiency-focused code generation metric Beyond and establish a benchmark to evaluate leading Code LLMs using this metric in Section 3.
@@
  • Baselines. In Section 4, we detail our extensive analysis of two baselines to enhance code efficiency while maintaining functional correctness. Experiment results reveal that despite Code LLMs excelling in functional correctness, there is still considerable potential to elevate efficiency.
@@
Table 1: A comparison of Mercury to existing NL2Code benchmarks. Mercury distinguishes itself by including a set of distilled high-quality solutions and a dedicated test case generator for each task. * signifies that the solution number can be further expanded by the data collection framework.
@@
Mercury 256 Online + Filters $\infty$ 18.4 * 3
@@ 2 Mercury Datasets
We initiate this work by collecting public programming tasks on Leetcode [23]. Subjecting these questions to a series of filters, we distilled them down to 1,889 high-quality tasks. A difficulty-balanced subset of 256 tasks was randomly selected to form the Mercury-eval benchmark, which obtains an average of 18.4 solutions for each problem. The remaining tasks have been designated as the Mercury-train dataset for baseline training (detailed data distribution is listed in Appendix Table 6). To enhance clarity within this paper, we employ Mercury to denote Mercury-eval unless otherwise specified.
Data Schema. As illustrated in Figure 2, Mercury offers a unified data schema to streamline the evaluation procedure and bolster further development endeavors. The data scheme encompasses these principal components: (1) Task Description contains the task instruction interpreted into a plain text format, along with illustrative examples and constraints of inputs and outputs. (2) Test Case Generator refers to a Python code snippet designed to automatically produce a comprehensive set of test cases in accordance with the specifications laid out in the task description. (3) Solutions are sampled from Leetcode historical submissions. Each solution within Mercury has undergone rigorous testing, and Locality-Sensitive Hashing [17] is employed to prevent the inclusion of any identical solutions. (4) Prompts and Entry Points where prompts act as the initiating prefixes for LLM code generation and entry points denote the start point for code execution. We delineate the definition of Mercury fields in the Appendix Table 5.
Task Filters. Mercury tasks originate from public programming problems on Leetcode. To assure the quality and uniformity of the dataset, we distilled gathered tasks based on the following conditions:
  1. 1.
@@
    Restricted Data Structure. Above the inherent Python data types, Mercury also incorporates two custom data types: Binary Tree and Linked List (the specific structure definitions can be found in Appendix Figure 4), which increases Mercury’s diversity and escalates its difficulty level. Tasks that contain other data structures will be removed.
  3. 3.
@@
    Unique Outputs. Certain Leetcode tasks may permit non-unique answers. For example, a result list can be returned in any order. Evaluating all possible answers can drastically complicate the test case verification process. To eliminate this problem, we harness the corresponding test case generator to generate $N$ test cases $T_{i}=\langle Input_{i},Output_{i}\rangle\ s.t.\ i\in\{0,1,\cdots,N\}$ and execute $T$ on different solutions $S_{m}\ s.t.\ m\in\{0,1,\cdots,M\}$ to observe if all $Output_{i}=S_{m}(Input_{i})\ s.t.\ i\in\{0,1,\cdots,N\}$ remain identical. Any tasks that potentially yield non-unique answers were subsequently excluded.
@@
Task Difficulty. Most existing NL2Code benchmarks predominantly comprise simplistic tasks, leading to a situation where LLMs of varied capabilities address most tasks effortlessly and yield indistinguishable high scores [47, 18]. To alleviate this issue, Mercury inherits the difficulty categorization from Leetcode, i.e., Easy, Medium, and Hard. The stratification aims to probe the upper bounds of Code LLM capabilities, delivering a more evident distinction between various Code LLMs.
@@
Test Case Generator. Manual creation of test cases can be a laborious process. To gather sufficient test cases to conduct an exhaustive assessment, we assign a test case generator for each evaluation task, which can produce a full range of test cases to thoroughly evaluate the functional correctness and code efficiency of given solutions. Specifically, We feed $pretty\_content$ into GPT-4 [34] to generate an initial test case generator snippet. To confirm the effectiveness of the initial generator, we subsequently create 24 test cases by the generator and submit these cases to the Leetcode Online Judge (OJ) system. Should any of the generated test cases not pass the LeetCode OJ validation, we manually revise the generator until all generated cases can be successfully validated.
3 Code Efficiency Metric
@@
In the domain of software development, code efficiency can be defined as the absolute code runtime for executing a given test case set [7]. Nonetheless, a primary obstacle in benchmarking code efficiency is normalizing runtime measurements across disparate environments. For instance, a sub-optimal solution might have a faster absolute runtime on high-performance hardware than an optimal solution on low-performance hardware. Moreover, different operation systems and code interpreters may also fluctuate the code runtime. Therefore, absolute runtime fails as a consistent and reliable code efficiency benchmark metric. To address this issue, an intuitive approach involves modeling a devoted runtime distribution for each task and calculating the average runtime percentiles of LLM solution samples over the runtime distribution. With this idea in mind, we proposed a normalized code efficiency metric Beyond:
@@
Where $N$ is the total number of tasks, and $K$ denotes the size of LLM solution samples. For a specific task $n\in N$, $R^{n}$ is the runtime array corresponding to the collected historical solutions, and $r_{k}^{n}\ s.t.\ k\in K$ denotes the runtime for the $k$-th LLM solution. $clip$ is a function to constraint the value $r_{k}^{n}$ in the range $[min(R^{n}),max(R^{n})]$. Runtime is defined as the period from the solution instantiation to the evaluation across all test cases, culminating with a successful termination (More engineering details can be found in Appendix Section A.3). Since any case failure of the $k$-th solution results in $r_{k}^{n}\rightarrow+\infty$ and then $p_{k}^{n}=0$, Beyond can reflect functional correctness as well.
Untrusted Code Execution. Since most Code LLMs are trained on an extensive code corpus from unverified sources, there is an intrinsic risk that these models may produce malicious code when driven by specific meticulous prompts [8]. The direct execution of synthesized code raises significant security concerns. To alleviate the risk of running untrusted code, we engage a robust sandbox to execute code in an isolated environment. Sandbox details are deliberated in Appendix A.3
Environment-agnostic Evaluation. To ensure fair comparison across diverse configurations, we run each task $n$ with corresponding test cases locally and aggregate their runtimes into the runtime array $R^{n}$. Appendix Figure 10 illustrates the Beyond score of two LLMs (‘deepseek-coder-33b’ and ‘deepseek-coder-6.7b’) over three distinct hardware specifications: the micro-tier (0.25 CPU cores), the small-tier (0.5 CPU cores), and the standard-tier (1 CPU core). The results demonstrate that Beyond remains consistent over different hardware configurations.